Work Out Question
1. The demand for petrol rises from 500 to 600 Barrels when the price of a particular scooter is reduced from Birr. 25000 to Birr.22000. Find out the cross elasticity of demand for the two. What is the nature of their relationship?
A company has the following demand equation
Q= 1000–3000P+10A
Q = Quantity demanded
P = Product Price
A = Advertisement expenditure
Assume that P = 3 and A = 2000
2. Suppose the firm drops the price to Birr. 2.50 would this be beneficial.
3. Suppose the firm raises the price to Birr. 4.00 While increasing its advertisement expenditure by 100 would this be beneficial? Explain
1
Expert's answer
2020-05-28T11:16:25-0400

1. The cross elasticity of demand is:

"Ed = \frac {600 - 500} {22000 - 25000}×\frac{22000 + 25000} {600 + 500} = -1.42."

So, these goods are complements.

2. If the firm drops the price to Birr. 2.50, it wouldn't be beneficial, because the demand is inelastic.

3. This would be beneficial, because as the demand is inelastic, the total revenue would increase.